please can u send me the answers to these and the full balnce method of how you worked it out
1. 6a=48
2. 4y-15=21
3.5p+9=39
4.3(n+5=36
5.7(2n-3)=49

Answer by stanbon(75887)

please can u send me the answers to these and the full balnce method of how you worked it out
1. 6a=48
Divide both sides by 6 to get:
a = 8
--------------------
2. 4y-15=21
Add 15 to get: 4y = 36
Divide by 4 to get:
y = 9
-------------------
3.5p+9=39
Subtract 9 to get: 5p = 30
Divide by 5 to get:
p = 6
------------------------
4.3(n+5)=36
divide by 3 to get:
n+5 = 12
Subtract 5 to get:
n = 7
-----------------------
5.7(2n-3)=49
Divide by 7
2n-3 = 7
Add 3
2n = 10
Divide by 2
n = 5
===============
Cheers,
Stan H.
================